OpenStudy (anonymous):

y = -0.57x^2 + 6.79x - 16.68 how to put in general form showing all steps?

OpenStudy (dumbcow):

its in standard form...do you mean vertex form?

OpenStudy (anonymous):

i need to know how to put it in general form

OpenStudy (anonymous):

from standart to general

OpenStudy (dumbcow):

\[y = a(x-h)^{2} +k\] this form?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think its something with completing the square?

OpenStudy (dumbcow):

ok you have to use completing the square haha

OpenStudy (anonymous):

lol well i figured:p i just for got how to do that :/

OpenStudy (dumbcow):

step 1: move constant to other side, and factor out coefficient of x^2 term \[\rightarrow y+16.68 = -0.57(x^{2}-11.91x)\] step 2: take half of x term and square it, then multiply by -.57 and add it to other side \[\rightarrow y+16.68 -20.22 = -0.57(x^{2}-11.91x+35.475)\] step 3: factor perfect square, simplify and move constants back to right side \[\rightarrow y = -0.57(x-5.956)^{2}+3.54\]

OpenStudy (anonymous):

where did the 11.91x come from?

OpenStudy (dumbcow):

6.79/-.57 = -11.91..

OpenStudy (anonymous):

oh.. yeah i see my bad lol

OpenStudy (dumbcow):

it would be better to practice with easier numbers to get the hang of it first http://worksheetplace.com/mf/Completing-the-Square-1.pdf

OpenStudy (anonymous):

lol yeah im trying to figure out what you did but ill try the link thanks:)

OpenStudy (dumbcow):

sorry if i skipped steps...but i literally did exactly what i stated in each step

OpenStudy (anonymous):

http://www.khanacademy.org/math/algebra/quadtratics/v/completing-the-square BEST EXPLANATION EVER